OpenStudy (anonymous):

AM WILLING TO BECOME A FAN FOR WHO EVER WHO HELPS ME!!! HOW DO YOU FACTOR 5X^2-7X+2 ???

OpenStudy (anonymous):

5x^2 - 7x - 2 = 5x^2 - 5x - 2x - 2 = 5x(x - 1) - 2(x - 1) = (x -1)(5x - 2) Ans: The factors are (x - 1) and (5x - 2)

OpenStudy (anonymous):

oh whoops.. mine was for -2

OpenStudy (anonymous):

i think you can fix that tho

OpenStudy (anonymous):

thanks lots all